# Bohao Wu   Qingyun Wang   Yue Guo

### v1

@@ Abstract
Personalizing jargon detection and explanation is essential for making technical documents accessible to readers with diverse disciplinary backgrounds. However, tailoring models to individual users typically requires substantial annotation efforts and computational resources due to user-specific finetuning. To address this, we present a systematic study of personalized jargon detection, focusing on methods that are both efficient and scalable for real-world deployment. We explore two personalization strategies: (1) lightweight fine-tuning using Low-Rank Adaptation (LoRA) on open-source models, and (2) personalized prompting, which tailors model behavior at inference time without retaining. To reflect realistic constraints, we also investigate hybrid approaches that combine limited annotated data with unsupervised user background signals. Our personalized LoRA model outperforms GPT-4 by 21.4% in F1 score and exceeds the best performing oracle baseline by 8.3%. Remarkably, our method achieves comparable performance using only 10% of the annotated training data, demonstrating its practicality for resource-constrained settings. Our study offers the first work to systematically explore efficient, low-resource personalization of jargon detection using open-source language models, offering a practical path toward scalable, user-adaptive NLP system.
@@
Bohao Wu   Qingyun Wang   Yue Guo University of Illinois at Urbana-Champaign {bohaowu,qingyun4,yueg}@illinois.edu
@@ 1 Introduction
Figure 1: Bridging the disciplinary gap under constraints of limited compute and annotation: Our approach leverages PEFT, prompting, and mixed data to enable efficient and effective adaption.
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide spectrum of applications, showcasing sophisticated natural language understanding and generation ability Zhou et al. (2024); Wang et al. (2024); Baek et al. (2025). As research increasingly transcends disciplinary boundaries, LLMs are being leveraged to facilitate the integration of diverse fields, offering new avenues for processing complex, cross-disciplinary data and generating novel insights Leto et al. (2024); Ramoneda et al. (2024); Lu et al. (2024); Jiang et al. (2025).
However, a significant impediment to seamless interdisciplinary collaboration and knowledge sharing lies in the challenge of domain-specific jargon Barnett and Doubleday (2020). Researchers often encounter specialized terminology outside their core expertise, which can hinder effective communication and understanding across disciplines Han et al. (2018); Lucy et al. (2023). This is particularly salient when scholars need to engage with literature from unfamiliar domains, potentially leading to misinterpretations and impeding the synthesis of knowledgeGlasziou et al. (2020). Our primary motivation is to provide personalized assistance to researchers in navigating this linguistic landscape, enabling a more accessible and efficient engagement with interdisciplinary content. A crucial first step in realizing this goal is the accurate identification of domain-specific jargon tailored to the individual researcher’s background and expertise. Therefore, this paper focuses on personalized jargon identification.
Figure 2: Schematics of the supervised, unsupervised and the combined training strategies for personalized LLMs.
Additionally, the increasing demand for user-centric AI has spurred significant interest in personalized Large Language Models (LLMs). While various strategies have been proposed, each presents its own limitations. Benchmarks like LaMP (Language Model Personalization) (Salemi et al., 2024) underscore the importance of this area but also reveal the complexities in achieving effective personalization across diverse tasks. Techniques focusing on adapting the model itself, such as One PEFT Per User (OPPU) (Tan et al., 2024b) and Personalized Pieces (Per-Pcs) (Tan et al., 2024a), demonstrate promising results in encoding user preferences through parameter-efficient fine-tuning or lightweight plug-in modules. However, these methods typically rely on supervised data derived from explicit user interactions or preferences, which can be sparse, costly to obtain, and may not fully generalize to unseen scenarios or capture implicit user needs. Furthermore, while efforts like HLLM (Chen et al., 2024) explore personalization in specific downstream tasks like recommendation, the applicability of such task-specific architectures to the broader challenge of personalized language understanding remains an open question. The need for data-efficient and broadly applicable personalization, especially when explicit user feedback is scarce, motivates our exploration of unsupervised personalized PEFT for jargon identification.
Addressing the challenges at the intersection of interdisciplinary communication and personalized language models, Guo et al. (2024) pioneered the use of LLMs, employing GPT-4 OpenAI (2023) to gauge out-of-domain scholars’ familiarity with jargon. Building upon their innovative work, which faced computational costs and suboptimal performance, our research aims to overcome these limitations by developing more efficient and tailored solutions leveraging lightweight, open-source LLMs for individual researchers. Furthermore, recognizing the data-intensive nature and potential generalization limits of prevalent supervised personalized PEFT approaches, this paper investigates the potential of unsupervised personalized PEFT to enhance jargon identification. Our contributions are:
  • Personalized Language Model Finetuning: We introduce a PEFT approach with LoRA that significantly improves personalized jargon identification over prior work (Guo et al., 2024) with minimal resources.
  • Personalized Prompting Strategies: We demonstrate that tailored prompting based on researcher background significantly influences jargon identification across domains, offering efficient performance optimization.
  • Personalization through Unsupervised Background Integration: We use unsupervised researcher data with minimal labeled data (10%) to achieve performance comparable to fully annotated models.
@@ 2 Method
2.1 Prompting Setup
We retrieve a series of background contexts to guide our personalized jargon identification. Specifically, given a pair of the research abstract and the corresponding entity of interest extracted from it, we retrieve four types of background contextual information (details in App. Table 2): 1) Metadata: We retrieve the annotator’s metadata from Semantic Scholar API Lo et al. (2020), including the abstract’s domain, the annotator’s publication number, the annotator’s reference count, and the annotator’s definition of their research domain. 2) Profile: As illustrated in Fig 2, we synthesize the researcher profile using the baseline model, which describes the annotator’s background and research information based on his/her publications. 3) Nearest Annotator: We leverage BM25 (Trotman et al., 2014) as our retrieval function, with the collection of annotator profiles serving as our document library, which we term the profile network. This method retrieves the annotator most similar to each target within the profile network, using their familiarity labels as contextual information for the target. 4) Nearest Abstract: Similarly, we apply BM25 to all the abstracts in the dataset, which we refer to as the abstract network. Then, we retrieve the most similar abstract within this network and use the target annotator’s familiarity labels as the contextual information. 5) For comparison, we also add a vanilla prompting, which means there is no input for related data, so the model has no information about the annotator.
2.2 Supervised Model Training
Based on the findings of Tan et al. (2024b), which demonstrate that LoRA (Hu et al., 2022) consistently outperforms other parameter-efficient fine-tuning (PEFT) methods, we adopt LoRA for model adaptation. Our objective in this stage is to assess whether applying LoRA to the complete dataset can lead to performance improvements beyond those achieved in previous studies (Guo et al., 2024). In this supervised setting, we maintained a consistent prompting strategy for both training and inference, aligning with the specific design of each context enhancement method detailed in Table 3. Following the Alpaca template (Taori et al., 2023), each input includes a jargon entity along with its corresponding abstract, placed in the “Input”, while the binary familiarity label is placed in the “Response”.
2.3 Unsupervised + Personalized PEFT
To assess the performance of unsupervised PEFT in our task, we deliberately avoid using task-specific formats during fine-tuning to prevent contamination. Instead, we compile a dataset composed of the titles and abstracts from each annotator’s previously published papers. We then apply LoRA fine-tuning for a few epochs. In cases where an annotator lacks sufficient publication data for effective training, we supplement the dataset by randomly retrieving papers from their self-defined research subdomain. This setup simulates a real-world scenario in which a new researcher has little or no labeled data available. Our goal is to explore whether a personalized model can still be trained effectively under such constraints, using only domain-relevant but unlabeled content.
3 Experiments and Results
3.1 Supervised Finetuning and Familiarity Classification
Figure 3: Model performance in familiarity classification for supervised versus unsupervised models, with various fine-tuning epochs under different prompting strategies. Evaluation is done upon the validation dataset. Each node is a combination of $1$ personalized models, specifically finetuned on 11 corresponding annotators. Here, GPT-4 and Oracle demonstrate previous SOTA model performance, denoting prompt-based GPT-4 and Oracle classifiers.
In Fig. 3 (left), we present the results of fine-tuning our baseline model, Llama-3.1 8B Instruct 4bit Dubey et al. (2024), using the full annotation dataset over various training epochs. Across all prompting strategies, a rapid performance increase is observed within the initial 20 epochs, followed by a plateau. Vanilla prompting achieves the highest effective F1 score of 0.82. Metadata-enhanced and profile-enhanced prompting exhibit similar performance, saturating around 0.80. Conversely, the nearest annotator and nearest abstract prompting strategies show lower performance, leveling off at approximately 0.75. All evaluations were conducted on the validation dataset. For comparison, Figure 3 (left) shows that our LoRA-based models significantly outperform prior state-of-the-art results from Guo et al. (2024)) (GPT-4 context learning: F1 0.64; oracle classifier: F1 0.72), even with simpler prompting strategies, demonstrating the effectiveness of supervised fine-tuning with LoRA in learning semantic connections between jargon and abstracts.
3.2 Unsupervised Finetuning
The performance of our unsupervised fine-tuned models, evaluated on the validation dataset across varying epochs and prompting strategies, is illustrated in Fig. 3 (right). The results indicate an initial performance dip upon commencement of fine-tuning, followed by a subsequent improvement, reaching its peak at 50 training epochs. Notably, the nearest annotator and nearest abstract prompting strategies yielded nearly 100% mismatch rates due to the unsupervised models’ lack of explicit task format knowledge and the longer context lengths associated with these strategies. Therefore, their results are excluded from the figure. Overall, unsupervised fine-tuning demonstrates a modest improvement over the baseline pre-trained models but does not achieve the performance levels of state-of-the-art supervised methods.
Figure 4: Validation performance plots for models fine-tuned under different strategies. To enable a consistent comparison across strategies, each involving datasets of varying text lengths, we use training steps as the x-axis. This also provides an approximate indication of the computational cost associated with each fine-tuning approach.
3.3 Mixed Finetuning on Supervised and Unsupervised Data
This section investigates the extent to which our fine-tuned models learn semantic knowledge versus simply aligning to the task template and explores the utility of unsupervised data when labeled data is scarce. To assess semantic learning, we observed a significant decline in performance saturation when training data was reduced to 1% and 10% (Fig. 4), with 1% models underperforming unsupervised methods. This performance decline, especially the 1% model underperforming unsupervised methods, indicates that the supervised models have learnt from the semantic information within the annotated content, rather than merely aligning to the task template. To address the challenge of limited annotation, we explored a mixed strategy, finding that integrating unsupervised researcher background data with just 10% of the original training set achieved a comparable performance (up to 79.1%) to fully supervised models, significantly outperforming models trained solely on the 10% subset. This demonstrates that leveraging unsupervised data with minimal annotation can yield competitive results, reducing the burden of extensive manual labeling.
@@ 4 Conclusions
To address the barrier of domain-specific jargon in interdisciplinary communication, we presented a novel framework for personalized jargon identification. By adapting lightweight LLMs with LoRA, we significantly outperformed prior work with minimal resources. We also demonstrated the substantial impact of personalized prompting based on the researcher’s background. Specifically, our data-efficient technique leveraging unsupervised data achieved near fully supervised performance with minimal annotation. This work offers a resource-conscious and scalable approach to building personalized tools that lower barriers to interdisciplinary understanding.
@@
One limitation of our current work is the reliance on a specific dataset focused primarily on computer science researchers and a limited number of out-of-domain areas. While this allowed for a controlled evaluation of our personalized techniques, the generalizability of our findings to a broader range of interdisciplinary domains and diverse researcher backgrounds requires further investigation. Future work should explore the application and evaluation of our framework on more heterogeneous datasets encompassing a wider spectrum of academic disciplines and research profiles to assess its robustness and adaptability in more varied real-world scenarios.
@@
In this paper, we utilized anonymized data from a pre-existing dataset, raising ethical considerations regarding the privacy and responsible use of researcher background information in future implementations. We acknowledge the potential for our personalized models to inherit or amplify biases present in pre-trained models or training data, necessitating careful evaluation across diverse user groups to ensure equitable performance. Furthermore, we recognize the importance of clarifying jargon without oversimplification and the potential for over-reliance on such tools to impact researchers’ own interdisciplinary language development. Finally, we advocate for responsible development to prevent unintended consequences like the creation of echo chambers. Ongoing evaluation and community discussion are essential to navigate these ethical complexities.

### v2

@@ Abstract
Personalizing jargon detection and explanation is essential for making technical documents accessible to readers with diverse disciplinary backgrounds. However, tailoring models to individual users typically requires substantial annotation efforts and computational resources due to user-specific finetuning. To address this, we present a systematic study of personalized jargon detection, focusing on methods that are both efficient and scalable for real-world deployment. We explore two personalization strategies: (1) lightweight fine-tuning using Low-Rank Adaptation (LoRA) on open-source models, and (2) personalized prompting, which tailors model behavior at inference time without retaining. To reflect realistic constraints, we also investigate hybrid approaches that combine limited annotated data with unsupervised user background signals. Our personalized LoRA model outperforms GPT-4 by 21.4% in F1 score and exceeds the best performing oracle baseline by 8.3%. Remarkably, our method achieves comparable performance using only 10% of the annotated training data, demonstrating its practicality for resource-constrained settings. Our study offers the first work to systematically explore efficient, low-resource personalization of jargon detection using open-source language models, offering a practical path toward scalable, user-adaptive NLP system.  [^1] Code can be found at: anonymous GitHub repository.
@@
Bohao Wu   Qingyun Wang   Yue Guo University of Illinois at Urbana-Champaign   William & Mary {bohaowu,yueg}@illinois.edu  qwang16@wm.edu
@@ 1 Introduction
Figure 1: Bridging disciplinary gap with limited compute and annotation: Our approach leverages PEFT, background-aware prompting, and mixed supervision to enable efficient and effective model adaption.
Large Language Models (LLMs) are increasingly used to support interdisciplinary research by helping scholars navigate diverse and domain-specific texts Leto et al. (2024); Ramoneda et al. (2024); Lu et al. (2024); Jiang et al. (2025). However, a persistent barrier to effective interdisciplinary collaboration is the prevalence of domain-specific jargon Barnett and Doubleday (2020); Strober (2006). Researchers often struggle to interpret specialized terminology outside their core expertise, leading to miscommunication Han et al. (2018); Choi and Pak (2007), impaired knowledge integration Lucy et al. (2023), and ultimately slow scientific discovery Glasziou et al. (2020); Daniel et al. (2022); van Helden et al. (2024). While prior work has developed NLP methods to identify and simplify scholarly jargon using general-purpose corpora like Wikipedia as proxies for reader knowledge (Gardner and Davies, 2013; Tanaka-Ishii and Terada, 2011; Guo et al., 2022, 2021), these approaches remain limited by their lack of personalization. A researcher’s background significantly influences their familiarity with domain-specific terms (Gooding and Tragut, 2022; Guo et al., 2024), suggesting that individualized models could more effectively determine which terms require explanation.
To address this challenge, we focus on the task of personalized jargon identification: automatically detecting domain-specific terms that may be unfamiliar to an individual researcher based on their background. Our goal is to make interdisciplinary content more accessible by leveraging LLMs in a personalized, data-efficient, and scalable manner.
Figure 2: To enable personalized jargon detection, we first fine tune a LoRA based PEFT model using supervised, unsupervised, and hybrid training strategies that reflect real world scenarios with varying levels of annotation availability. Next, we enhance the contextual understanding of the target researcher through a range of background aware prompting methods, including vanilla, metadata based, profile enhanced, and two oracle prompting configurations using nearest neighbor techniques (researcher and abstract), to generate personalized familiarity predictions.
Recent efforts in personalized language models, such as LaMP (Salemi et al., 2024), OPPU (Tan et al., 2024b), and Per-Pcs (Tan et al., 2024a), have shown promise by adapting models to user preferences via parameter-efficient fine-tuning (PEFT). However, these methods often rely on costly supervised data or explicit annotation, limiting generalizability. While systems like HLLM (Chen et al., 2024) target personalization task, they do not directly address the broader challenge of personalized language understanding in scholarly context.
Our work builds on Guo et al. (2024), who introduced the first benchmark with scholar annotations for interdisciplinary jargon and explored the use of GPT-4 OpenAI (2023) for personalized jargon detection. While innovative, their method relies on computationally expensive LLM prompting and requires access to detailed metadata and annotated examples, limiting scalability, generalization, and applicability in low-resource or cold-start settings. We propose a more efficient and generalizable solution using open-source LLMs and unsupervised PEFT techniques. Our contributions are:
  • Lightweight Personalized PEFT: Introducing the first LoRA-based PEFT method for personalized jargon detection that outperforms prior work while significantly reducing computational overhead.
  • Background-Aware Prompting: Designing personalized prompting strategies using researcher background, yielding efficient performance gains across domains.
  • Unsupervised Personalization: Integrating unsupervised user data achieves strong performance with only 10% labeled supervision, enhancing accessibility and generalizability.
@@ 2 Method
To model each researcher’s background, we first retrieve contextual information §2.2 and explore three fine-tuning settings: supervised, unsupervised, and hybrid §2.3.
2.1 Experimental Setup
We use the personalized jargon detection dataset from Guo et al. (2024), which contains 11k term familiarity (i.e., familiar or unfamiliar) annotations provided by 11 computer science researchers for terms extracted from 100 paper abstracts. To our knowledge, it is the only interdisciplinary jargon dataset with high-quality annotations, rich personal metadata, and accompanying published papers from the annotators. We follow the same data split as Guo et al. (2024), dividing the dataset into 60/20/20 for the train, validation, and test sets. We selected Llama-3.1 8B Instruct 4bit Dubey et al. (2024) as our baseline model for personalized jargon detection based on its strong instruction-following ability, low mismatch rate, and competitive performance in preliminary evaluations across several state-of-the-art LLMs. We evaluate model performance using the effective F-1 Score, which adjusts for output format errors by penalizing predictions with a high mismatch rate. Details in App. §A.
2.2 Background-Aware Prompting
Building on prior work Guo et al. (2024); Tan et al. (2024b), we design a set of prompting strategies that incorporate varying levels of researcher-specific context into the model input. These strategies differ in both the type and richness of the background information provided:
  • Metadata: Structured features including the annotator’s self-defined subfield (e.g., NLP, computer vision), publication count, average references, year of first publication, and the domain of the current abstract. These serve as lightweight indicators of expertise and familiarity.
  • Profile-enhancement: Following prior work on user modeling in personalized recommendation Tan et al. (2024b), we use the baseline model to generate a natural language summary of the annotator’s research background based on their metadata.
To explore upper-bound performance, we also include two oracle prompting configurations based on social recommendation Kang et al. (2022); Guy et al. (2009); Goldberg et al. (1992):
  • Nearest Annotator: We use BM25 Trotman et al. (2014) to retrieve the most similar annotator based on profile text, and use their familiarity labels for the most similar terms as proxy input.
  • Nearest Abstract: We retrieve the most similar abstract using BM25 and use the target annotator’s familiarity labels for its terms as context.
As a baseline, we include a vanilla prompting setup in which no contextual information about the annotator is provided.
2.3 Lightweight Personalized PEFT
To ensure consistency across different personalization settings, we adopt the Alpaca format (Taori et al., 2023), which is widely used in instruction fine-tuning. Each training instance is structured into three components: an Instruction, an Input (target abstract and term), and a Response (binary familiarity label). We use this standardized format to unify model interaction across different training strategies. Task-specific instructions and prompt examples are shown in Table 3. We evaluate three training strategies with varying supervision levels:
  • Supervised: We adopt LoRA (Hu et al., 2022) for PEFT, following findings from Tan et al. (2024b) demonstrating its strong performance. In this setting, the Instruction field includes background context retrieved as described in §$2.$
  • Unsupervised: To simulate low-resource scenarios, we fine-tune using only unlabeled data, including titles and abstracts from each annotator’s previous publications. For annotators with limited publication history, we supplement the data with papers from their self-defined subdomain. This reflects practical cases where personalized models must be trained on domain-relevant but unlabeled content.
  • Hybrid: To examine the trade-off between annotation effort and personalization quality, we also explore a hybrid approach that combines a small amount of labeled data with the unlabeled corpus. This setting aims to inform future annotation strategies by evaluating how much supervision is sufficient for reasonable model performance.
3 Results
Figure 3: Validation performance of supervised (a), unsupervised (b), and hybrid (c) models in jargon familiarity detection. Each point represents a model fine tuned for one of the 11 annotators using various prompting strategies. GPT-4 and Oracle baselines indicate prior state-of-the-art performance from Guo et al. (2024). For hybrid models, training steps are used on the x-axis to enable consistent comparison across datasets of different sizes and to reflect relative computational cost.
Supervised fine-tuning outperforms GPT-4 with strong gains across prompting strategies
Fig. 3 (a) shows the validation performance of our supervised model trained on the full annotated dataset across different prompting strategies and training epochs. All methods show rapid performance gains in the first 20 epochs, after which improvements plateau. Among prompting strategies, vanilla prompting achieves the highest F1 score of 0.82, suggesting that even without explicit background context, the model benefits significantly from full supervision. Metadata-enhanced and profile-enhanced prompts yield comparable results, both saturating around 0.80, indicating modest additional benefit from lightweight personalization. In contrast, oracle style prompts (i.e., nearest annotator and nearest abstract) achieve lower performance, plateauing around 0.75. This may reflect noise introduced during retrieval or label transfer from similar but non-identical contexts, despite their oracle style design. On the test set (Table 1), we use 20 and 50 epochs for training the supervised and unsupervised models, respectively. Our supervised LoRA-tuned models substantially outperform the prior state-of-the-art reported by Guo et al. (2024), where GPT-4 with context learning achieved an F1 score of 0.64 and the oracle classifier reached 0.72. Notably, even our models using vanilla prompting exceed these benchmarks. These results highlight the effectiveness of full supervision combined with parameter-efficient fine-tuning in capturing the semantic relationships between research abstracts and jargon familiarity.
Models F-1 Score $\uparrow$ Recall Precision
Best results from Guo et al. (2024)
Oracle Classifier 71.9±1.7 76.0±2.1 68.2±2.1
GPT-4 64.2±1.5 98.7±0.5 47.6±1.6
Supervised Training
Vanilla 77.9±1.2 77.8±2.2 78.0±0.5
Matadata-enhanced 76.8±1.1 76.1±2.9 77.7±1.0
Profile-enhanced 76.6±1.0 73.7±2.0 79.9±1.6
Nearest Annotator 72.1±3.0 70.3±4.1 74.5±2.7
Nearest Abstract 77.8±1.1 78.3±1.7 77.5±2.5
Unsupervised Training
Published Papers 54.6±5.1 77.5±7.6 45.0±0.8
Hybrid Training
1% Sup 53.5±2.8 56.0±4.5 51.5±3.2
10% Sup 63.6±2.9 59.8±5.1 69.0±1.7
10% Sup + Unsup 77.0±1.1 78.9±2.6 75.4±0.7
Table 1: Performance of fine-tuned models on the test set. Unless specified, results use vanilla prompting. The $\pm$ values represent standard deviations across 3 repeated runs to illustrate consistency.
Unsupervised fine-tuning without annotations shows limited effectiveness
In the unsupervised setting, models were fine-tuned solely on each annotator’s published papers, without familiarity annotations. Although performance improves slightly over time, overall results remain poor, confirming that publication history alone fails to capture familiarity judgments, consistent with prior findings Haghani (2023). We exclude the nearest annotator and nearest abstract settings from this experiment, as they require access to annotated familiarity labels and are therefore not applicable in the unsupervised scenario.
Unsupervised personalization enables efficient adaptation with minimal supervision.
Our findings show that integrating unsupervised user publication data substantially boosts performance when combined with just 10% of labeled training data, achieving an F1 score of 77.0. This mixed strategy not only approaches the performance of fully supervised models but also significantly outperforms models trained on the same limited labeled data alone. These results underscore the potential of leveraging unlabeled, domain-relevant data to reduce annotation costs while preserving personalization quality. This approach improves the scalability and accessibility of personalized NLP systems, particularly in settings where manual annotation is expensive or infeasible. Additional qualitative analyses, such as model generalization to related tasks, performance across individual annotators, and domain-specific behavior, are in App.§B.
@@ 4 Conclusions
In this work, we addressed the challenge of domain-specific jargon as a barrier to interdisciplinary communication by introducing a novel framework for personalized jargon identification. Leveraging lightweight large language models fine-tuned with LoRA, our approach achieves significant performance gains over prior work while maintaining computational efficiency. We demonstrated the effectiveness of incorporating personalized prompts grounded in a researcher’s background, and showed that integrating unsupervised user data with minimal labeled supervision can nearly match fully supervised performance. Together, these contributions offer a scalable and resource-efficient path toward building personalized NLP tools that enhance accessibility and foster cross-disciplinary understanding.
@@
One limitation of our current work is its reliance on a specific dataset, which is primarily focused on computer science researchers and encompasses a limited number of out-of-domain areas. While this allowed for a controlled evaluation of our personalized techniques, the generalizability of our findings to a broader range of interdisciplinary domains and diverse researcher backgrounds requires further investigation. Future work should explore the application and evaluation of our framework on more heterogeneous datasets that encompass a wider spectrum of academic disciplines and research profiles, to assess its robustness and adaptability in more varied real-world scenarios.
@@
In this paper, we utilized anonymized data from a pre-existing dataset, raising ethical considerations regarding the privacy and responsible use of researcher background information in future implementations. We acknowledge the potential for our personalized models to inherit or amplify biases present in pre-trained models or training data, necessitating careful evaluation across diverse user groups to ensure equitable performance. Furthermore, we recognize the importance of clarifying jargon without oversimplification and the potential for over-reliance on such tools to impact researchers’ own interdisciplinary language development. Finally, we advocate for responsible development to prevent unintended consequences like the creation of echo chambers. Ongoing evaluation and community discussion are essential for navigating these ethical complexities.